Help me solve this. Everytime I do a check, it does not balance. 5x-8x+10x=28-10x+3x
Answer by stanbon(75887)

5x-8x+10x=28-10x+3x
7x = 28-7x
14x = 28
x = 2
==========
Cheers,
Stan H.